Phathu needs R10 500 in ten months' time to buy herself a new lens for her camera. Two months ago she deposited R9 000 in a savings account at a simple interest rate of 11,5% per year. How mu h money will Phathu still need to buy the lens ten months from now?

1.     R465,00

2.     R229,50

3.     R637,50

4.     R408,67

5.     None of the above


1
Expert's answer
2021-05-13T08:35:38-0400

Amount required after 10 months = R10,500

Two months ago amount deposited = R 9000

Interest Rate =11.5%

Interest rate per month (r)"= 11.5\%/12 = 0.958333 = 0.00958333"

Months of compounding for 9000 is 10+2 (n)= 12

Therefore, amount after 10 months "= 9000\times(1+r)^n"

"=9000\times(1+0.00958333)^{12}"

"=9000\times[1.12125928]"

"=10091.33"

Therefore, Amount required "=10500 - 10091.33 = 408.67"

The correct answer is  4.     R 408.67